ab=2/9ba, where a + b = 9
Answer by AnlytcPhil(1806)

 ab = 2/9ba, where a + b = 9


We cannot tell whether that first equation is
       2
ab = —————, which you type as ab = 2/(9ba)
      9ba


      2
ab = ——— a, which you type as ab = [2/(9b)]a
      9b

or

      2
ab = ——— ba, which you type as (2/9)(ba)
      9

Please repost

Edwin
AnlytcPhil@aol.com